# GAI_arXiv.bib

### v2

@@ IV Conclusion
This study examines whether collective reasoning among generative agents can facilitate novel and coherent thinking that leads to innovation. To achieve this, it proposes GAI, a new LLM-empowered framework designed for reflection and interaction among multiple generative agents to replicate the process of analogy-driven innovation. The framework’s functionality is evaluated using Dyson’s invention of the bladeless household fan as a case study, assessing the extent to which the core ideas of the innovation can be replicated through a set of fictional technical documents.
The experimental results demonstrate that models with internal states significantly outperformed those without, achieving higher average scores and lower variance. Notably, the model with five heterogeneous agents equipped with internal states successfully replicated the key ideas underlying the Dyson Air Multiplier. This indicates that the internal state module enables agents to refine their ideas, resulting in the construction and sharing of more coherent and comprehensive concepts.
-A Format of user’s current thought
a. Functional Similarlities
The functional similarities between the target domain and the ejector:
- In both domains, ...
The points in agent_2’s opinion on the topic of the functional similarities that I agree with or feel interesting are as follows:
The points in agent_2’s opinion on the topic of the functional similarities that contain contradictions or missing information are as follows:
b. Mechanical Differences
How the mechanisms in the ejector that produce the common functionalities differ from those in the target domain:
c. Solutions
How the functionalities or mechanisms of the ejector and the target domain should be modified or retained to solve the challenges of the target domain:
d. New Challenges
The potential new challenges arising from applying the ejector’s mechanisms:
e. New Opportunities
The additional opportunities generated by applying the ejector’s mechanisms:
-B Sample prompts for dialogue
Phase A: Functional similarities
Hello. The field of the product we are developing is referred to as the ’target domain,’ and the field of reference products for development is referred to as the ’source domain.’ The target domain for this session is household fan. The source domain is ejector. Read the following documents about the target domain and the source domain, and provide your thoughts on the core functional similarities between the two domains, starting with what you believe to be the most essential similarities. However, do not discuss the differences between the two domains, the challenges within each domain, their possible solutions yet. Please respond within 170 words.
Phase B: Mechanical differences
Thank you. Next, please share your thoughts on how the mechanisms in the source domain that produce the common functionalities differ from those in the target domain, starting with the most significant differences. However, do not discuss the challenges within each domain or their solutions yet. Please respond within 170 words.
Phase C: Solutions
Thank you. Now, let’s consider how applying the discussed differences in mechanisms to the target domain could help solve its challenges. Please share your thoughts on what aspects of the source domain’s mechanisms should be adapted and how they should be integrated into the target domain to address the challenges in the target domain. Additionally, explain what functionalities or mechanisms of the target domain should be retained and what should be discarded, in order of importance, starting with the most critical points. Please respond within 170 words.
Phase D: New challenges
Thank you. Now, let’s consider whether applying the source domain’s mechanisms to the target domain in the manner discussed could lead to any unforeseen significant challenges. If you believe such challenges might arise, please share your thoughts, listing them in order of severity. If you do not believe any major challenges would arise, please state that as your answer. Please respond within 170 words.
Phase E: New opportunities
Thank you. Now, let’s consider the opposite perspective. Could applying the source domain’s mechanisms to the target domain create additional opportunities beyond merely solving the target domain’s existing challenges? If you believe such new opportunities exist, please share your thoughts, listing them in order of importance. If you do not believe any additional opportunities would be generated, please state that as your answer. Please respond within 170 words.
[…]

### v3

@@ 5 CONCLUSIONS
This study examines whether collective reasoning among generative agents can facilitate novel and coherent thinking that leads to innovation. To explore this, it proposes a new LLM-empowered framework designed to enable reflection and interaction among agents, replicating analogy-driven innovation. It demonstrates, using Dyson’s bladeless fan as a case study, that an architecture recursively processing agents’ internal states, combined with a dialogue scheme designed for analogy-driven innovation, enables the construction and sharing of coherent and comprehensive concepts.
Future research directions include: (1) examining agents’ behavior under more complex organizational structures, (2) utilizing external tools such as CAD and integrating multimodal capabilities into generative agents, and (3) demonstrating collaboration with human teams through partnerships with real-world companies.